Circular No. 8290

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2004Z IN MCG +10-19-85 Further to IAUC 8259, T. Boles reports his discovery of an apparent

supernova (mag 18.5) on unfiltered CCD images taken on Feb. 19.119 and 19.790 UT with a 0.35-m reflector. The new object is located at α = 13^{h}32^{m}17^{s}.38, δ = +60^{o}23′43′′.4 (equinox 2000.0), which is ≈ 8′′.3 east and 5′′.8 north of the center of MCG +10-19-85. SN 2004Z is not present on Boles’ images from 2003 May 25 and Apr. 2 (limiting mag 19.5) or on Digitized Sky Survey red (1997) or blue (1995) plates.

IRAS 05436–0007 G. Masi, “Tor Vergata”, University of Rome, reports that CCD im-

ages (scale 0′′.87/pixel), acquired remotely by F. Mallia and himself with the Tenagra II telescope near Nogales, AZ, yield the following R-band magnitudes (in a 6′′ aperture) for the star in outburst at the southern end of the nebula (cf. IAUC 8284, 8289): Feb. 11.16 UT, 16.5; 14.15, 16.4; 18.14, 16.9. Further to IAUC 8287, K. Hornoch adds that this same star brightened in R by ∼ 0.20 mag between Feb. 11.913 and 12.908. Masi adds that the nebula itself was present on tricolor CCD images obtained by R. Croman on 2003 Dec. 30 (with the nebula possibly a bit fainter) and by M. Halderman on 2003 Nov. 26 (with the nebula clearly fainter than now), but a similar image taken by Halderman on 2003 Feb. 6 shows no nebula at this location. First and second Palomar Sky Survey red and blue images, together with Kitt Peak images taken in 1998 and 2000, show a faint nebulosity corresponding to the northern part of the currently visible nebula (around the location of HH 22). Masi and D. di Cicco (Sky and Telescope) further note that there is a nebular region visible, with a similar shape to the IRAS 05436–0007 nebula now, on an image taken by E. Kreimer on 1966 Oct. 22 on Tri-X film (see Sky Tel. 39, 27). Di Cicco adds that a search of the Sky and Telescope photo archive suggests that the nebula was invisible on numerous dates from the late 1970s through early 2003.

SUPERNOVA 2004R Corrigendum. The right ascension of SN 2004R (cf. IAUC 8281) should

read 3^{h}32^{m}41^{s}.30 (not 31^{s}.30).

2004 February 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green